OpenStudy (anonymous):

Factor the GCF and then factor the resulting binomial: 50x^3-128x

OpenStudy (anonymous):

x(50x^2-128) 2x(25x^2-64) 2x(5x-8)(5x+8) x=-8/5 x=0 x=8/5